Mickley Companys plantwide predetermined overhead rate is $19.00 per direct labor-hour and its direct labor wage rate hour.
Answer #1

Solution 1:

Direct labor hours for Job A-500 = Direct labor cost / $14 per hour = $280 / $14 = 20 hours

Overhead Assigned to Job A-500 = Direct labor hours * $19 per DLH = 20 * $19 = $380

Total manufacturing cost = Direct materials + direct labor + Overhead assigned = $260 + $280 + $380 = $920

Solution 2:

Unit product cost = Total manufacturing costs / units = $920 / 50 = $18.40 per unit